OpenStudy (anonymous):

translation with the same effect as (x+7, y-2) followed by (x-3, y+2)

OpenStudy (cwrw238):

+7 - 3 = +4 -2 + 2 = 0

OpenStudy (anonymous):

okay thank you. How would I do that myself though? Like what would I do to get that answer.

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

is it just x-x over y-y

OpenStudy (amistre64):

start with x, add 7, subtract 3 start with y, subtract 2, add 2

OpenStudy (amistre64):

it tells you exactly what to do in the transformations

OpenStudy (aum):

(x+7, y-2) followed by (x-3, y+2) What does the transformation (x+7, y-2) mean? It says add 7 to the x-coordinate and subtract 2 from the y-coordinate. If, for example, we start with the point (2,3) and apply the transformation (x+7, y-2) the point would be transformed to become (9,1). The second transformation, (x-3, y+2), says subtract 3 from the x-coordinate and add 2 to the y-coordinate. If this transformation is applied to (9, 1) it will become (6, 3). Now if we combine both transformations: (x+7, y-2) followed by (x-3, y+2) then add 7 to the x-coordinate and subtract 2 from the y-coordinate and then subtract 3 from the x-coordinate and add to the y-coordinate. So to the x-coordinate we add 7 and subtract 3. In effect we add +7 - 3 = +4. To the y-coordinate we subtract 2 and add 2. Effectively it is: -2 + 2 = 0. Therefore, the combined transformations is: (x+4, y). If we apply the transformation (x+4, y) to the earlier example of (2,3) it becomes (6,3) which is what we got earlier by applying the two transformations one step at a time.

OpenStudy (anonymous):

@aum Thank you so much for explaining it

OpenStudy (anonymous):

@amistre64 thank you

OpenStudy (amistre64):

youre welcome :) aums explanation is more indepth of course ;)

OpenStudy (anonymous):

@amistre64 they both helped a lot though. :)

OpenStudy (aum):

You are welcome. It is nice that you wanted to learn than just getting the answer.

OpenStudy (anonymous):

@aum Yeah, I would rather be able to learn then just have someone give me the answer. I want to be able to do it myself and get it correct when I have a to take test on it instead of having to get it wrong. lol